the line that contains (1,7) and has the same y-intercept as the 2x+3y=6
Answer by lwsshak3(11628)

the line that contains (1,7) and has the same y-intercept as the 2x+3y=6
**
Standard form of equation for a straight line: y=mx+b, m=slope, b= y-intercept
2x+3y=6
3y=-2x+6
y=-2x/3+2
y-intercept=2
..
equation of line with the same y-intercept:
y=mx+2
solve for m using coordinates of given point (1,7)
7=m*1+2
m=5
equation of line: y=5x+2